OpenStudy (anonymous):

solving quadratic equations. z(z-1)=0

jimthompson5910 (jim_thompson5910):

If z(z-1)=0 then z = 0 or z-1=0

jimthompson5910 (jim_thompson5910):

Keep going to completely solve for z

OpenStudy (anonymous):

i don't understand how to do the problem? could you explain please? sorry. haven't been in algebra for quite sometime, but the help is much appreciated!

jimthompson5910 (jim_thompson5910):

If two numbers multiply to 0, then one of them must be 0

OpenStudy (anonymous):

The quadratic EQUATION is \[solution =\frac{b \pm \sqrt{b^2-4 a c}}{2 a}\] which holds true for any \[a z^2 + b z + c = 0\] Where a, b, and c are real numbers and z is complex (and, therefore, can be real). You have to get z(z-1) = 0 in the form I showed above to employ the quadratic equation. Then again, if you are given z(z-1) as is, in factored form, then it is easier the other way...

jimthompson5910 (jim_thompson5910):

So if z(z-1)=0 then either z is zero or z-1 is zero, which means z=0 or z-1=0 which becomes z=0 or z=1

OpenStudy (anonymous):

oh my goodness, that makes perfect sense now! thank you so much @jim_thompson5910!

jimthompson5910 (jim_thompson5910):

you're very welcome

jimthompson5910 (jim_thompson5910):

I'm glad it's making sense now

OpenStudy (anonymous):

would you mind helping with another problem and explaining as well? if that wouldn't be too much of a bother? y^2+y-30=0 i've gotten to this far in the equation: y^2=0 or y-30=0

jimthompson5910 (jim_thompson5910):

not quite

jimthompson5910 (jim_thompson5910):

you can only say that if y^2 times y-30 was equal to 0

jimthompson5910 (jim_thompson5910):

so you have to factor the left side first

jimthompson5910 (jim_thompson5910):

Do you know how to factor this?

OpenStudy (anonymous):

i'm factoring it right now

OpenStudy (anonymous):

well you don't have to factor, if you read what I wrote

OpenStudy (anonymous):

@jim_thompson5910 okay, i worked it like this: y^2+y-30=0 (y+5)(y-6)=0 y+5=0 or y-6=0 y=-5 or y=6 is that correct?

jimthompson5910 (jim_thompson5910):

close, but not quite

jimthompson5910 (jim_thompson5910):

-6 + 5 = -1, but we want the two numbers to add to positive 1

jimthompson5910 (jim_thompson5910):

but you have the right idea (minus that mistake)

OpenStudy (anonymous):

so it'd be: y^2+y-30=0 (y+6)(y-5)=0 y+6=0 or y-5=0 y=-6 or y=5 correct?

jimthompson5910 (jim_thompson5910):

much better

OpenStudy (anonymous):

thank you! could you help with this problem as well? lol 3c^2=5c

jimthompson5910 (jim_thompson5910):

get everything to one side

jimthompson5910 (jim_thompson5910):

what do you get when you do that

OpenStudy (anonymous):

i simplified it wrong, i got a fraction..

jimthompson5910 (jim_thompson5910):

that's normal (and in this case correct)

jimthompson5910 (jim_thompson5910):

depends on which fraction of course

OpenStudy (anonymous):

5/3 or 1and 2/3

jimthompson5910 (jim_thompson5910):

that's one solution

jimthompson5910 (jim_thompson5910):

what's the other

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

@jim_thompson5910 still helpin me?

jimthompson5910 (jim_thompson5910):

oh sry, i got distracted

jimthompson5910 (jim_thompson5910):

no, 2 isn't the other solution

OpenStudy (anonymous):

well darn, idk then

jimthompson5910 (jim_thompson5910):

3c^2=5c 3c^2-5c = 0 c(3c - 5) = 0 c = 0 or 3c-5=0 c = 0 or c = 5/3 So the two solutions are c = 0 or c = 5/3

jimthompson5910 (jim_thompson5910):

of course, you can write 5/3 as 1 2/3 (a mixed number)

jimthompson5910 (jim_thompson5910):

not sure if you did it this way, but this is the way I recommend doing it

OpenStudy (anonymous):

yeah i definitely did it some crazy way, but that way makes more sense

jimthompson5910 (jim_thompson5910):

alright I'm glad it makes sense

OpenStudy (anonymous):

okay what about this problem... 10x^2+9x+2=0 (-5x+2)(-2x+1)=0 -5x+2=0 or -2x+1=0 idk what to do after that^^

jimthompson5910 (jim_thompson5910):

unfortunately 10x^2+9x+2 doesn't factor to (-5x+2)(-2x+1)

OpenStudy (anonymous):

oh okay.

jimthompson5910 (jim_thompson5910):

try it again and tell me what you get

OpenStudy (anonymous):

umm i guessed but i got (-5x+3)(3x+1)

jimthompson5910 (jim_thompson5910):

still no 10*2 = 20 (first and last multiply to 20) Notice how 5+4 = 20 and 5*4 = 20 So break up 9x into 5x+4x and factor by grouping 10x^2+9x+2 10x^2+5x+4x+2 (10x^2+5x)+(4x+2) 5x(2x+1)+(4x+2) 5x(2x+1)+2(2x+1) (5x+2)(2x+1) So 10x^2+9x+2 factors to (5x+2)(2x+1)

OpenStudy (anonymous):

so then it would be 5x+2 =0 or 2x+1=0 right?

jimthompson5910 (jim_thompson5910):

yes, keep going

OpenStudy (anonymous):

5-2=0 or 2-1=0?

jimthompson5910 (jim_thompson5910):

not sure what you're doing there...

OpenStudy (anonymous):

wait x=5-2 or x=2-, i messed up

jimthompson5910 (jim_thompson5910):

closer, but a bit off

jimthompson5910 (jim_thompson5910):

5x+2 =0 5x+2-2 =0-2 5x =-2 5x/5 = -2/5 x = - 2/5

jimthompson5910 (jim_thompson5910):

2x+1=0 2x+1-1=0-1 2x = -1 2x/2 = -1/2 x = -1/2

OpenStudy (anonymous):

okay thank you

jimthompson5910 (jim_thompson5910):

yw